# Sparse Hanson-Wright Inequalities with Applications

### v2

@@ 2 Previous results and comparison
Addressing the asymptotic behavior, [12] established the limiting distributions for the random quadratic form $S_{n}=\sum_{i,j}a_{ij}x_{i}x_{j}$, where $A=(a_{ij})$ is a symmetric, $\{0,1\}$-valued, diagonal-free matrix and $x_{i}$ are i.i.d. Bernoulli random variables with parameter $p$. Their study focused on the sparse regime where $\mathbb{E}(S_{n})=p^{2}\sum_{i,j}a_{ij}=O(1)$.
We conclude the literature review with two open questions. It is unclear to us if a sparsity-adaptive Hanson-Wright inequality exists for Bernoulli random vectors that is optimal for all sparsity regimes. Taking $p_{i}$ to be a constant independent of $n$ in (2.6) and (2.7) does not recover the Hanson-Wright inequality for sub-gaussian random vectors in [35]. And (2.6) and (2.7) yields different bounds depending on various norms of $A$. For sparse sub-gaussian random vectors, it would be interesting to unify our Theorem 1 and Zhou’s inequality in [47].
@@ 3.1 Sparse Wigner matrices: local law and eigenvector delocalization
where $\widetilde{H}_{ij}$ ($1\leq i\leq j\leq n$) are independent centered random variables. Each entry $\widetilde{H}_{ij}$ has the form $\widetilde{H}_{ij}=x_{ij}\xi_{ij}$, where $x_{ij}$ and $\xi_{ij}$ are independent, with $x_{ij}\sim\mathrm{Ber}(p)$ and $\xi_{ij}$ having mean 0, variance 1, and being $\alpha$-subexponential. We assume that for any $r\geq 1$, $\max_{i,j}\|\xi_{ij}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$, where $K\geq 1$ is a constant. Here $p=p(n)\in(0,1)$.
Denote the eigenvalues of $H$ as $\lambda_{1}(H),\cdots,\lambda_{n}(H)$ and their corresponding normalized eigenvectors as $u_{1},\cdots,u_{n}$. When $np\to\infty$, it is known that the empirical spectral distribution (ESD) of $H$, defined as $\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\lambda_{i}(H)}$, converges weakly to the semicirle law in probability [44, 41]. When $\xi_{ij}=1$, $H$ is the adjacency matrix of Erdős-Rényi graph, and its spectral properties, including the local law and eigenvector delocalization [41, 24], universality [18, 19, 25], and extreme eigenvalues [9, 5], have received considerable attention in recent years. When $\xi_{ij}$ is a bounded random variable, several extensions have been obtained in [5, 40, 6, 17]. When $p=\frac{c}{n}$, large deviation principle of the spectral edge was studied in [20].
We establish the local law and complete eigenvector delocalization for $H$, analogous to He, Knowles, and Marcozzi’s work [24] that studied the (normalized) adjacency matrix of the Erdős-Rényi random graph $G(n,p)$ in the supercritical regime $np\geq C\log n$. Our methodology follows the framework of [24], a method that was developed through a series of works [19, 18] (see references therein) and is comprehensively surveyed in [10]. The key innovation in our approach lies in the application of the large deviation inequalities developed in this paper to handle the sparse $\alpha$-subexponential setting.
@@
Consider the matrix $H$ in (3.1) and its Green function $G(z)$. For any $D>0$ and $\delta\in(0,1)$, there exists $L\equiv L(\delta,D)$ such that the following holds. Let $p\geq LC_{\alpha}^{2}K^{4}\frac{(\log n)^{\max\{2,\frac{2}{\alpha}\}}}{n}$. For $z=E+\mathrm{i}\eta\in\mathbf{S}$ with $\eta\geq LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, we have
@@
A standard consequence of the local law is the semicircle law on small scales given below. Recall that $\mu_{n}$ is the ESD of $H$.
@@
Under the assumption of Theorem 8, for any $D>0$ and interval $I\subset[-5,5]$ satisfying $|I|\geq LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, we have
@@
Another consequence of the local law is the complete delocalization of the (normalized) eigenvectors $\{u_{i}\}_{i=1}^{n}$ of $H$.
THEOREM 10 (Complete eigenvector delocalization).
Consider the matrix $H$ in (3.1). For any $D>0$, there exist $L=L(D)$ and $n_{0}=n_{0}(D)$ such that for $p\geq LC_{\alpha}^{2}K^{4}\frac{(\log n)^{1+\max\{\frac{2}{\alpha},1\}}}{n}$ and $n\geq n_{0}$, we have
@@
The study of the complete eigenvector delocalization has a rich history (see survey [33]). For the adjacency matrix of the Erdős-Rényi random graph $G(n,p)$, [24] showed that all eigenvectors satisfy $\|u_{i}\|_{\infty}\leq\frac{n^{(\log n)^{-1/2}}}{\sqrt{n}}$ with high probability when $p\geq C\frac{\sqrt{\log n}}{\sqrt{n}}$, improving upon the bounds established in [41]. The conjectured optimal bound is $\|u_{i}\|_{\infty}\lesssim\frac{\sqrt{\log n}}{\sqrt{n}}$, which has been proven for bulk eigenvectors of Wigner matrices with bounded entries [42]. For all eigenvectors, the optimal delocalization for Wigner matrices with sub-exponential entries was established in [11]. Our bound in Theorem 10 appears to be $\sqrt{\log n}$ away from this optimal bound in the dense case, yet represents the current best result for Wigner matrices with sparse $\alpha$-subexponential entries. The proof of Theorem 10 follows by combining two key ingredients: the local law and an estimate of the spectral norm of $H$. The complete details are presented in Appendix C.
By the same proof method with the sparse Hanson-Wright inequality for Bernoulli random vectors in (2.3), we can obtain $\|u_{i}\|_{\infty}\lesssim\frac{\log n}{\sqrt{n}}$ for $p\geq\frac{C\log^{2}n}{\sqrt{n}}$, which improves the delocalization estimate of [24] in the super-critical regime $p\geq\frac{C\log^{2}n}{\sqrt{n}}$.
@@
The results presented in this section can be extended in several directions. First, the matrix $H$ can be complex-valued, requiring only cosmetic modifications to our proofs. The entries may satisfy more general moment conditions as discussed in Remark 1.1. Moreover, the diagonal entries can have a common variance that differs from the variance of the off-diagonal entries. A more substantial generalization involves sparse Wigner-type matrices with heterogeneous sparsity parameters $p_{ij}$, as studied in [17]. Such extensions can be achieved using the quadratic vector equation developed in [4]. While these generalizations are natural and tractable, we do not pursue their details in this work.
@@ 3.2 Concentration of the Euclidean norm of sparse random vectors
Another application of our sparse Hanson-Wright inequality concerns the estimation of the Euclidean norm of random vectors, an application previously considered in [35, 22] in the non-sparse setting. For a sparse $\alpha$-subexponential random vector $X$ and a deterministic matrix $B$, we establish the following concentration result for the squared norm $\|BX\|^{2}$.
@@
Consider $X=(X_{1},\dots,X_{n})^{T}$ with $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ is a mean-zero, unit-variance random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$ for all $r\geq 1$ and some $\alpha>0$. Let $B\in\mathbb{R}^{m\times n}$ be an deterministic matrix with columns $B_{1},\dots,B_{n}$. For any $s>0$, with probability at least $1-C^{\prime}\exp(-cs)$, the following holds:
@@
$+sK^{2}\max_{1\leq i\leq n}\Big{\{}p\sum_{j\neq i}|\langle B_{i},B_{j}\rangle|,\|B_{i}\|^{2}\Big{\}}+s^{\max\{\frac{2}{\alpha},2\}}K^{2}\|B^{T}\|_{2,\infty}^{2}.$ (3.6)
Note that in the above result, $\|B^{T}\|_{2,\infty}=\max_{i}\|B_{i}\|$ denotes the maximum Euclidean norm among all columns of $B$.
Let us compare our result with the existing bound in [22] for the special case $\alpha=1$, although both their result and ours apply to $\alpha\in(0,1]$. Consider a sparse sub-exponential random vector $X$ with i.i.d. entries $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ is a mean-zero, unit-variance random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$. These conditions imply $\|X_{i}\|_{\Psi_{1}}\leq CK$ for a universal constant $C>0$. According to [22, Corollary 2.2], with probability at least $1-C^{\prime}\exp(-cs)$,
@@
4 Proof of Theorem 1
The key input of the proof of Theorem 1 is the following moment estimates, whose proof is deferred to Section 4.1.

### v3

@@ 2 Previous results and comparison
Addressing the asymptotic behavior, [16] established the limiting distributions for the random quadratic form $S_{n}=\sum_{i,j}a_{ij}x_{i}x_{j}$ with symmetric, $\{0,1\}$-valued, diagonal-free $A$ and i.i.d. Bernoulli $x_{i}$, focusing on the sparse regime where $\mathbb{E}(S_{n})=O(1)$.
Comparison to Theorem 1 and Theorem 5
We now compare our results with these key findings from the literature discussed above. For $\alpha\in[1,\infty]$, the bound (2.4) aligns with our Theorem 1, while our Theorem 5 provides an improvement over the term $\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{1/2}$ in the large deviation regime. The result in (2) represents the special case where all sparse parameters $p_{i}=1$. Notably, (2.3), which represents the sparse bounded case, is consistent with (2) when $\alpha=1$, while incorporating the sparse parameters $p_{i}$.
For $\alpha\in(0,1)$, our Theorem 1, when all $p_{i}=1$, yields a weaker bound than (2). This can be seen from the corresponding moment inequalities:
$r^{\frac{2+\alpha}{2\alpha}}K^{2}\|A\|_{2,\infty}\leq r^{\frac{1}{\alpha}+\frac{1}{2}}K^{2}\|A\|_{1,\infty}\|A\|_{\max}\leq r^{\frac{2}{\alpha}}K^{2}\|A\|_{\max}+rK^{2}\|A\|_{1,\infty}.$ (2.5)
Using a similar argument, when $X_{i}$’s are centered Bernoulli random variables, (2.3) outperforms our Theorem 1 when $\alpha=\infty$. For $\alpha=2$, unlike (2.2), taking $p_{i}=1$ in either Theorem 1 or Theorem 5 does not recover the Hanson-Wright inequality for sub-gaussian random vectors in [52]. However, in Theorem 5 (see also Theorem 3), we improve the large deviation term to $\left(\frac{t}{\|A\|_{\max}}\right)^{1/2}\log^{1-\frac{1}{\alpha}}\left(\frac{t\|A\|_{\max}}{C_{\alpha}\Lambda^{2}}\right)$ for $t$ sufficiently large. These additional log factors are essential for our application for local law and eigenvector delocalization of sparse random matrices in Section 3.1.
A natural direction for future work is to derive a refined concentration inequality that unifies our results (Theorem 1 for $\alpha\in(0,1]$ and Theorem 5 for $\alpha\in(1,\infty]$) and the above results. Our techniques naturally extend to polynomials of independent random variables, as studied in [32, 31, 54, 53]; these extensions will appear in forthcoming work.
Linear forms of random vectors
The concentration of the sum of independent random variables $\sum_{i=1}^{n}a_{i}X_{i}$ has been an area of intense research focus, with a rich body of literature. Our main contribution in this direction is the establishment of new Bennett-type inequalities in Theorem 4 (see also Theorem 3) for the sum of sparse $\alpha$-subexponential random variables when $\alpha>1$. We will briefly review some key results.
For bounded random variables, the classical Bennett’s inequality [15] (see [17, Theorem 2.9]) provides the following bound: For $\max_{i}X_{i}\leq K$,
$\mathbb{P}\left(\sum_{i=1}^{n}a_{i}(X_{i}-\mathbb{E}(X_{i}))\geq t\right)\leq\exp\left(-\frac{\nu^{2}}{K^{2}\|a\|_{\infty}^{2}}h\left(\frac{K\|a\|_{\infty}t}{\nu^{2}}\right)\right),$ (2.6)
where $\nu^{2}=\sum_{i=1}^{n}a_{i}^{2}\mathbb{E}(X_{i}^{2})$ and $h(u)=(1+u)\log(1+u)-u$ for $u>0$. In particular, when $t\gtrsim\frac{\nu^{2}}{K\|a\|_{\infty}}$, the tail bound on the RHS of (2.6) is approximately $\exp\left(-\frac{t}{CK\|a\|_{\infty}}\log\left(\frac{K\|a\|_{\infty}t}{\nu^{2}}\right)\right).$ Compared to our Theorem 3 where $\alpha=\infty$, since $\nu^{2}\leq K^{2}\sum_{i}a_{i}^{2}p_{i}=K^{2}\lambda^{2}$, replacing $\nu^{2}$ with $K^{2}\lambda^{2}$ in this tail bound yields exactly (1.6) up to constants. As shown in [48, Example 2.4], the large deviation behavior captured by Bennett’s inequality achieves optimal sharpness (up to universal constants) for bounded random variables. Improved and refined versions of Bennett’s inequality for the sum of independent bounded random variables have been developed, incorporating the Lambert W function; see, for instance, [66], [36], and [45].
Although an exact analogue of our Theorem 2 for sparse $\alpha$-subexponential random variables has not been found, a similar conclusion might be achievable through existing moment estimates combined with appropriate conditioning techniques. In this direction, we highlight two relevant results from [40]. In [40, Theorem 3.1], the authors leverage results from [42] to establish
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{i}X_{i}\right|\geq t\right)\leq 2\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{2}\|a\|_{2}^{2}},\left(\frac{t}{KL_{\alpha}(a)}\right)^{\alpha}\right\}\right),$
where $L_{\alpha}(a)=\|a\|_{\infty}$ if $\alpha\leq 1$ and $L_{\alpha}(a)=\|a\|_{\frac{\alpha}{\alpha-1}}$ if $\alpha>1$. The bounds above are sharp with respect to their dependence on the vector $a$. However, these results do not capture the dependence on the sparsity parameters $p_{i}$. The tail bounds that scale with variance are presented in [40, Theorem 3.2], utilizing proof techniques from [1]:
$2\exp\left(-\min\left\{C\frac{t^{2}}{\sum_{i}a_{i}^{2}p_{i}\mathbb{E}\xi_{i}^{2}},C_{\alpha}\left(\frac{t}{K\|a\|_{\infty}\log^{\frac{1}{\alpha}}(n+1)}\right)^{\min\{\alpha,1\}}\right\}\right).$ (2.7)
The factor $(\log(n+1))^{1/\alpha}$ is shown to be sharp by [1, Section 2.2], assuming only $\|X_{i}\|_{\psi_{\alpha}}<\infty$ and variance-based norm bounds. In contrast, our Theorem 2 eliminates this logarithmic term in the subexponential tail by employing a variance proxy $K^{2}\sum_{i}a_{i}^{2}p_{i}$ in the sub-gaussian tail.
@@ 3.1 Sparse Wigner matrices: local law and eigenvector delocalization
where $\widetilde{H}_{ij}$ ($1\leq i\leq j\leq n$) are independent centered random variables with the following assumptions:
ASSUMPTION 1.
Each entry $\widetilde{H}_{ij}$ has the form $\widetilde{H}_{ij}=x_{ij}\xi_{ij}$, where $x_{ij}$ and $\xi_{ij}$ are independent, with $x_{ij}\sim\mathrm{Ber}(p)$ and $\xi_{ij}$ having mean 0, variance 1. Here $p=p(n)\in(0,1)$. We assume that for any $r\geq 1$, $\max_{i,j}\|\xi_{ij}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$, where $K\geq 1$ is a constant. When $\alpha=\infty$, this is interpreted as the $\xi_{ij}$’s being bounded by $K$.
Denote the eigenvalues of $H$ as $\lambda_{1}(H),\cdots,\lambda_{n}(H)$ and their corresponding $\ell_{2}$-normalized eigenvectors as $u_{1},\cdots,u_{n}$. When $np\to\infty$, it is known that the empirical spectral distribution (ESD) of $H$, defined as $\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\lambda_{i}(H)}$, converges weakly to the semicirle law in probability [62, 58]. When $\xi_{ij}=1$, $H$ is the adjacency matrix of Erdős-Rényi graph, and its spectral properties, including the local law and eigenvector delocalization [58, 34], universality [27, 28, 35], and extreme eigenvalues [12, 5], have received considerable attention in recent years. When $\xi_{ij}$ is a bounded random variable, several extensions have been obtained in [5, 57, 8, 26]. When $p=\frac{c}{n}$, the large deviation principle of the spectral edge was studied in [29, 30].
We establish the local law and complete eigenvector delocalization for $H$, analogous to He, Knowles, and Marcozzi’s work [34] that studied the (normalized) adjacency matrix of the Erdős-Rényi random graph $G(n,p)$ in the supercritical regime $np\geq C\log n$. As shown in [6], this is the optimal sparsity level (up to a constant factor) for complete eigenvector delocalization. To the best of our knowledge, this is the first local law and complete delocalization result for sparse $\alpha$-subexponeitial random matrices down to the near-optimal sparsity $p\geq\frac{\mathrm{polylog}(n)}{n}$ and for unbounded sparse sub-gaussian random matrices down to the optimal sparsity $p\geq\frac{C\log n}{n}.$
Our methodology follows the framework of [34], a method that was developed through a series of works [28, 27] (see references therein) and is comprehensively surveyed in [13]. The key innovation in our approach lies in the application of the large deviation inequalities developed in this paper to handle the sparse $\alpha$-subexponential setting.
@@
Consider the matrix $H$ in (3.1) under Assumption 1 and its Green function $G(z)$. For any $D>0$ and $\delta\in(0,1)$, there exists $L\equiv L(\delta,D)$ such that the following holds. Let $p\geq LC_{\alpha}^{2}K^{4}\frac{(\log n)^{\max\{2,\frac{2}{\alpha}\}}}{n}$. For $z=E+\mathrm{i}\eta\in\mathbf{S}$ with $\eta\geq LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, we have
@@
A standard consequence of the local law is the semicircle law on small scales, as given below. Recall that $\mu_{n}$ is the ESD of $H$.
@@
Under the assumption of Theorem 6, for any $D>0$ and interval $I$ satisfying $|I|\geq LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, we have
@@
Another consequence of the local law is the complete delocalization of the (normalized) eigenvectors $\{u_{i}\}_{i=1}^{n}$ of $H$. Note that compared to Theorem  6, we require a slightly stronger lower bound on $p$ due to a technical condition on the spectral norm concentration of $H$ (see Lemma B.1).
THEOREM 8 (Complete eigenvector delocalization for sparse Wigner matrix with $\alpha$-subexponential entries).
Consider the matrix $H$ in (3.1) under Assumption 1. For any $D>0$, there exist $L=L(D)$ and $n_{0}=n_{0}(D)$ such that for $p\geq LC_{\alpha}^{2}K^{4}\frac{(\log n)^{\max\{1+\frac{2}{\alpha},2\}}}{n}$ and $n\geq n_{0}$, we have
@@
The study of the complete eigenvector delocalization has a rich history (see survey [49]). For the adjacency matrix of the Erdős-Rényi random graph $G(n,p)$, [34] showed that all eigenvectors satisfy $\|u_{i}\|_{\infty}\leq\frac{n^{(\log n)^{-1/2}}}{\sqrt{n}}$ with high probability when $p\geq C\frac{\log n}{n}$, improving upon the bounds established in [58]. The conjectured optimal bound is $\|u_{i}\|_{\infty}\lesssim\frac{\sqrt{\log n}}{\sqrt{n}}$, which has been proven for bulk eigenvectors of Wigner matrices with bounded entries [60]. For all eigenvectors, the optimal delocalization for Wigner matrices with subexponential entries was established in [14]. Our bound in Theorem 8 appears to be $\sqrt{\log n}$ away from this optimal bound in the dense case, yet represents the current best result for Wigner matrices with sparse $\alpha$-subexponential entries. The proof of Theorem 8 follows by combining two key ingredients: the local law and an estimate of the spectral norm of $H$. The complete details are presented in Appendix B.1.
The conclusion of Theorem 8 can be extended to non-first eigenvectors of the adjacency matrix of the Erdős-Rényi random graph $G(n,p)$ with minor modifications to the proof, following the approach of [34]. When $\alpha=\infty$, our result improves upon [34] by tightening the upper bound on the infinity norms from $\frac{n^{(\log n)^{-1/2}}}{\sqrt{n}}$ to $\frac{\log n}{\sqrt{n}}$. However, Theorem 8 requires a more restrictive condition of $p\geq\frac{C\log^{2}n}{n}$, compared to their super-critical regime of $p\geq\frac{C\log n}{n}$.
In the proof of Theorem 8, we have adapted Theorem 2 and Theorem 1. By incorporating the improved bounds from Theorem 4 and Theorem 5, we can establish results analogous to [34] that hold up to the super-critical regime $p\gtrsim\frac{\log n}{n}$. Moreover, we demonstrate that these results extend to random matrices with sparse $\alpha$-subexponential entries, where $\alpha\in[2,\infty]$, in the super-critical regime. Note that $\beta=1-\frac{1}{\alpha}\in[\frac{1}{2},1].$ And when $\alpha\geq 2$, $H$ is a sparse sub-Gaussian random matrix.
THEOREM 9 (Complete eigenvector delocalization for sparse sub-Gaussian random matrices).
Consider the matrix $H$ in (3.1) with a parameter $\alpha\in[2,\infty]$. For any $D>0$ and $\varepsilon\in(0,1)$, there exist $L=L(D,\varepsilon)$ and $n_{0}=n_{0}(D,\alpha,K,\varepsilon)$ such that for $p\geq LC_{\alpha}^{2}K^{4}\frac{\log n}{n}$ and $n\geq n_{0}$, we have
$\mathbb{P}\left(\forall i:\|u_{i}\|_{\infty}\leq n^{-\frac{1}{2}+\varepsilon(\log n)^{-1+\frac{1}{2\beta}}}\right)\geq 1-2n^{-D}.$
To the best of our knowledge, this is the first complete eigenvector delocalization result for sparse sub-gaussian random matrices in the supercritical regime $p\geq\frac{C\log n}{n}$ beyond the bounded random variable case [58, 28, 34, 26]. A local law for $H$ when $\alpha\geq 2$ is also proved and we provide the detailed statement in Theorem 12. Sparse Wigner matrices correspond to sparse random graphs with random weights, which has gained significant interest recently in the large deviation theory of random matrices [30, 29, 9, 8]. For $\alpha=\infty$, corresponding to sparse bounded entries in $H$, the upper bound $\|u_{i}\|_{\infty}\leq n^{-\frac{1}{2}+(\log n)^{-\frac{1}{2}}}$ aligns with the result in [34]. The proof of Theorem 9 follows the same approach as Theorem 8, but incorporates the improved concentration inequalities for $\alpha\geq 2$, namely Theorem 4 and Theorem 5. A sketch of this proof is provided in Appendix B.2.
@@
The results presented in this section can be extended in several directions. First, the matrix $H$ can be complex-valued, requiring only cosmetic modifications to our proofs. Moreover, the diagonal entries can have a common variance that differs from the variance of the off-diagonal entries. A more substantial generalization involves sparse Wigner-type matrices with heterogeneous sparsity parameters $p_{ij}$, as studied in [26]. Such extensions can be achieved using the quadratic vector equation developed in [3]. While these generalizations are natural and tractable, we do not pursue their details in this work.
@@ 3.2 Concentration of the Euclidean norm of sparse random vectors
Another application of our sparse Hanson-Wright inequality concerns the estimation of the Euclidean norm of random vectors, an application previously considered in [52, 32] in the non-sparse setting. For a sparse $\alpha$-subexponential random vector $X$ and a deterministic matrix $B$, we establish the following concentration result for the squared norm $\|BX\|^{2}$ as a corollary of Theorem 1.
@@
Consider $X=(X_{1},\dots,X_{n})^{T}$ a random vector with independent centered sparse $\alpha$-subexponential components, where each $X_{i}$ satisfies $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr^{\frac{1}{\alpha}})^{r}$ for all $r\geq 1$ and some $\alpha\in(0,1]$. Let $B\in\mathbb{R}^{m\times n}$ be an deterministic matrix with columns $B_{1},\dots,B_{n}$. For any $s>0$, with probability at least $1-C^{\prime}\exp(-cs)$, the following holds:
@@
$+sK^{2}\max_{1\leq i\leq n}\Big{\{}p\sum_{j\neq i}|\langle B_{i},B_{j}\rangle|,\|B_{i}\|^{2}\Big{\}}+s^{\frac{2}{\alpha}}K^{2}\|B^{T}\|_{2,\infty}^{2}.$ (3.6)
Note that in the above result, $\|B^{T}\|_{2,\infty}=\max_{i}\|B_{i}\|$ denotes the maximum Euclidean norm among all columns of $B$. Theorem 10 is a direct consequence of Theorem 1. For $\alpha>1$, analogous results can be derived using Theorem 5; we omit the details here.
Let us compare our result with the existing bound in [32] for the special case $\alpha=1$, although both their result and ours apply to $\alpha\in(0,1]$. Consider a sparse subexponential random vector $X$ with i.i.d. entries $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ is a mean-zero, unit-variance random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$. These conditions imply $\|X_{i}\|_{\Psi_{1}}\leq CK$ for a universal constant $C>0$. According to [32, Corollary 2.2], with probability at least $1-C^{\prime}\exp(-cs)$,
@@
4 Generalization to general matrices and non-centered random vectors
Finally, we present two generalizations that combine the tail bounds for linear and quadratic forms. The first generalization is the Hanson-Wright inequality for a general matrix $A$. The second generalization is the Hanson-Wright inequality for sparse random variables with a diagonal-free matrix $A$, where we relax the assumption of centered random variables. For brevity, we only consider the special case of $\alpha=1$, though these results can be extended for arbitrary $\alpha>0$ using the results we present in Section 1.
COROLLARY 1.
Let $X=(X_{1},\cdots,X_{n})^{T}$ be a random vector whose components are independent centered sparse subexponential random variables, i.e., $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr)^{r}$ for all $r\geq 1$. Let $A$ be an $n\times n$ matrix. Then for any $t>0$,
$\leq$ $e^{2}\exp\left(-C\min\left\{\frac{t^{2}}{K^{4}(\sum_{i}a_{ii}^{2}p_{i}+\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j})},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\frac{1}{2}}\right\}\right),$
where
Note that the definition of $\gamma_{1,\infty}$ in (4.1) is consistent with the diagonal-free case in (1.4). Next, we provide the Hanson-Wright inequality for non-centered random variables.
COROLLARY 2.
Let $Y=(Y_{1},\ldots,Y_{n})^{T}$ be a random vector with independent sparse subexponential entries, where each $Y_{i}$ satisfies $\mathbb{E}(|Y_{i}|^{r})\leq p_{i}(Kr)^{r}$ for all $r\geq 1$. Let $A$ be an $n\times n$ diagonal-free matrix. Then there exists an absolute constant $C>0$ such that for any $t>0$,
$\mathbb{P}\left(\left|Y^{T}AY-\mathbb{E}(Y^{T}AY)\right|\geq t\right)$ (4.2)
$\leq e^{2}\exp\left(-C\min\left\{\frac{t^{2}}{K^{4}\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}+K^{2}\sum_{i}b_{i}^{2}p_{i}},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\frac{1}{2}}\right\}\right),$ (4.3)
where $b_{i}:=\max\{|\sum_{j\neq i}a_{ij}\mathbb{E}(Y_{j})|,|\sum_{j\neq i}a_{ji}\mathbb{E}(Y_{j})|\}$ and $\gamma_{1,\infty}$ is defined in (1.4).
5 Proofs of Theorem 1 and Theorem 2
The key input of the proof of Theorem 1 is the following moment estimates, whose proof represents the primary technical contribution of this work.